payton:

City A and City B had two different temperatures on a particular day. On that day, four times the temperature of City A was 8° C more than three times the temperature of City B. The temperature of City A minus twice the temperature of City B was −3° C. The following system of equations models this scenario: 4x = 8 + 3y x − 2y = −3 What was the temperature of City A and City B on that day? City A was 5° C, and City B was 4° C. City A was 3° C, and City B was −1° C. City A was 8° C, and City B was −3° C. City A was 5° C, and City B was −5° C.

carmelle:

@ramen

carmelle:

Did u even read the lesson?

ramen:

its ight i have all a's

jimthompson5910:

Solve the second equation for x x − 2y = −3 x − 2y+2y = −3+2y ..... adding 2y to both sides x = 2y-3 Then plug this into the first equation 4x = 8 + 3y 4(x) = 8 + 3y 4(2y-3) = 8+3y ... replace x with 2y-3 I'll let you solve for y

ramen:

Try graphing this equation.

ramen:

type the equation from the problem into the graphing calculator

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ramen https://www.desmos.com/calculator \(\color{#0cbb34}{\text{End of Quote}}\) use this

payton:

i still dont know what im doing

ramen:

ok i show you graph

payton:

okay

ramen:

1 attachment
ramen:

where do the 2 equations cross?

ramen:

what coordinate?

payton:

x=4 y=5

ramen:

yep and which of those answer choices have those numbers?

payton:

a

payton:

okay

carmelle:

yassssss

ramen:

smarty pants

carmelle:

yup